OpenStudy (anonymous):

Please help =( 2/x greater than or equal to 1/3

OpenStudy (imstuck):

you're solving this for x, right?

OpenStudy (anonymous):

\[\frac{ 2 }{ x }\ge \frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

yes i am

OpenStudy (imstuck):